# Sublinear Data Structures for Nearest Neighbor in Ultra High Dimensions

### v1

@@ 5.1.1 Upper Bound for Estimate of Distance between Query and its Nearest Neighbor
$\mathbb{E}[\|r_{i}-u\|_{1}]=\mathbb{E}\left[\sum_{b\in I}\left|r_{i}^{(b)}-u^{(b)}\right|\right]=\mathbb{E}\left[\sum_{t\in[T],b\in[d]}\mathbbm{1}_{\{\text{in iteration $t$, we add $b$ to $I$\}}}\left|r_{i}^{(b)}-u^{(b)}\right|\right]$
$=\sum_{t\in[T],b\in[d]}p^{(b)}\left|r_{i}^{(b)}-u^{(b)}\right|=\sum_{t\in[T],b\in[d]}p^{(b)}\left|\frac{c_{i}^{(b)}}{p^{(b)}}-\frac{q^{(b)}}{p^{(b)}}\right|=\sum_{t\in[T],b\in[d]}\left|c_{i}^{(b)}-q^{(b)}\right|=T\|c_{i}-q\|_{1}.$
@@
$\|r_{i^{*}}-u\|_{1}\leq\frac{1}{\delta^{\prime}}\mathbb{E}[\|r_{i}-u\|_{1}]\leq\frac{T}{\delta^{\prime}}\|c_{i^{*}}-q\|_{1}.\qed$
5.1.2 Coordinate Selection Preserves Distance Ratio
Definition 5.4 (Good region and bad region).
Consider two points $c_{1},c_{2}\in\mathbb{R}^{d}$, and a query point $q\in\mathbb{R}^{d}$. We define the good region w.r.t. $(c_{1},c_{2},q)$ as
$\mathsf{Good}(c_{1},c_{2},q)=\{z\in[d]\mid\min\{c_{1}^{(z)},c_{2}^{(z)}\}-\frac{8}{\epsilon\delta^{\prime}}|c_{1}^{(z)}-c_{2}^{(z)}|\leq q^{(z)}\leq\max\{c_{1}^{(z)},c_{2}^{(z)}\}+\frac{8}{\epsilon\delta^{\prime}}|c_{1}^{(z)}-c_{2}^{(z)}|\}$
Definition 5.5 (Restricted distance).
Lemma 5.6.
Proof.
@@ 5.1.2 Coordinate Selection Preserves Distance Ratio
$\|b-q\|_{1}+\|a-q\|_{1}\geq\mathsf{dist}_{\mathsf{Bad}(a,b,q)}(b,q)+\mathsf{dist}_{\mathsf{Bad}(a,b,q)}(a,q)\geq\sum_{z\in\mathsf{Bad}(a,b,q)}\left(\frac{8}{\epsilon\delta^{\prime}}+\frac{8}{\epsilon\delta^{\prime}}\right)|b^{(z)}-a^{(z)}|.$
@@
$\frac{\mathsf{dist}_{\mathsf{Good}(a,b,q)}(b,a)}{\|b-a\|_{1}}\geq\frac{14-\epsilon\delta^{\prime}}{16}\geq\frac{13}{16}.\qed$
Lemma 5.7.
  1. 1.
  2. 2.
  3. 3.
Proof.
@@
Notice that $\forall t\in[T],\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]=\mathsf{dist}_{\mathsf{Good}}(c_{i},q)$ and $\mathbb{E}[\sum_{z\in\mathsf{Good}}Q_{t,z}]=\mathsf{dist}_{\mathsf{Good}}(c_{j},q)$. Because
$\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]=\sum_{z\in\mathsf{Good}}\mathbb{E}[P_{t,z}]=\sum_{z\in\mathsf{Good}}p^{(z)}\cdot\frac{|q^{(z)}-c_{i}^{(z)}|}{p^{(z)}}=\sum_{z\in\mathsf{Good}}|q^{(z)}-c_{i}^{(z)}|=\mathsf{dist}_{\mathsf{Good}}(c_{i},q).$
@@
$\Pr\left[\left|\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}-\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right|\geq\epsilon_{2}\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right]\leq 2\exp\left(-\frac{\epsilon_{2}^{2}T\mathbb{E}[\sum_{z\in[d]}P_{t,z}]}{3(1+\frac{8}{\epsilon\delta^{\prime}})\|c_{i}-c_{j}\|_{1}}\right)$
$=2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{8}{\epsilon\delta^{\prime}})}\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},q)}{\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{j})}\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{j})}{\|c_{i}-c_{j}\|_{1}}\right)\leq 2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{8}{\epsilon\delta^{\prime}})}\cdot\frac{1}{2}\cdot\frac{13}{16}\right)\leq\frac{\delta^{\prime}}{2n},$
@@
$\Pr\left[\left|\sum_{t\in[T],z\in\mathsf{Good}}Q_{t,z}-\mathbb{E}[\sum_{t\in[T],z\in\mathsf{Good}}Q_{t,z}]\right|>\frac{T}{8}\cdot\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{j})\right]<2\exp\left(-\frac{T\cdot\frac{1}{8}\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{j})}{3(1+\frac{8}{\epsilon\delta^{\prime}})\|c_{i}-c_{j}\|_{1}}\right)$
$\leq 2\exp\left(-\frac{T}{24(1+\frac{8}{\epsilon\delta^{\prime}})}\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{j})}{\|c_{i}-c_{j}\|_{1}}\right)\leq\frac{\delta^{\prime}}{2n}.$
@@
The third item of the lemma is proved. ∎
Lemma 5.8.
Proof.
We know that $\|r_{i^{*}}-u\|_{1}\leq\frac{T}{\delta^{\prime}}\|c_{i^{*}}-q\|_{1}$ holds with probability $1-\delta^{\prime}$ by 5.3. Thus, it is sufficient to show the statement of the lemma holds under this assumption with probability $1-\delta^{\prime}$.
@@
Given these conditions, the following 5.9 gives a lower bound on the ratio $\frac{g_{1}^{\prime}+b_{1}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}$. We defer the proof of the claim to the appendix.
Claim 5.9.
@@
Applying 5.9, we get that $\frac{g_{1}^{\prime}+b_{1}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq 1+\epsilon\delta^{\prime}/4$, which implies $\frac{\|r_{i}-u\|_{1}}{\|r_{i^{*}}-u\|_{1}}\geq 1+\epsilon\delta^{\prime}/4$. By union bound, this inequality holds for all $i\in[n]$ with probability $1-\delta^{\prime}$. ∎
5.1.3 $\ell_{1}$ Dimension Reduction Preserves the Distance Ratio
Theorem 5.10 (Theorem 1 in [Ind06]).
Claim 5.11.
$\frac{\|c_{i}-q\|_{1}}{\|c_{i^{*}}-q\|_{1}}\leq 1+\epsilon$
with probability $1-3\delta^{\prime}$.
Proof.
Applying Theorem 5.10 with $m=O(\log(n/\delta^{\prime})/(\epsilon^{2}\delta^{\prime 2}))$, given any $v\in R$, it holds with probability $1-\delta^{\prime}/(n+1)$, that $F(Mv,Mu)\in[1-\epsilon\delta^{\prime}/100,1+\epsilon\delta^{\prime}/100]\|v-u\|_{1}$. By union bound, with probability $1-\delta^{\prime}$, it holds for all $v\in R$. By Lemma 5.8 it holds with probability $1-2\delta^{\prime}$ that for all $i\in[n]$, if $\frac{\|c_{i}-q\|_{1}}{\|c_{i^{*}}-q\|_{1}}\geq 1+\epsilon$, then $\frac{\|r_{i}-u\|_{1}}{\|r_{i^{*}}-u\|_{1}}\geq 1+\epsilon\delta^{\prime}/4$, thus $\frac{F(Mr_{i},Mu)}{F(Mr_{i^{*}},Mu)}>1$ with probability $1-3\delta^{\prime}$, i.e. the algorithm will not output $i$ as the index of the approximate nearest neighbor. ∎
5.1.4 Space and Query Complexity
Lemma 5.12.
$1\leq\sum_{b\in[d]}p^{(b)}\leq n$.
Proof.
Claim 5.13.
Proof.
Notice that $f(x)$ is a non-decreasing function, thus $f(x)\geq f(y)$, which proves $x-y\geq x^{\prime}-y^{\prime}$. Notice that $0\leq f(x)\leq\tau_{2}-\tau_{1}$, thus $f(x)\leq f(y)+(\tau_{2}-\tau_{1})$, which proves $x^{\prime}-y^{\prime}\geq x-y-(\tau_{2}-\tau_{1})$. ∎
Claim 5.14.
Proof.
Next, to prove (ii), first define $p_{b}=\max_{c_{i},c_{j}\in C,i\neq j}\frac{|c_{i}^{(b)}-c_{j}^{(b)}|}{\|c_{i}-c_{j}\|_{1}}$ and $p_{b}^{\prime}=\max_{c^{\prime}_{i},c^{\prime}_{j}\in C^{\prime},i\neq j}\frac{|{c^{\prime}_{i}}^{(b)}-{c^{\prime}_{j}}^{(b)}|}{\|c^{\prime}_{i}-c^{\prime}_{j}\|_{1}}$. We prove that $p_{b}-p_{b}^{\prime}\leq\frac{|c_{i_{0}}^{b}-c_{j_{0}}^{b}|}{\|c_{i_{0}}-c_{j_{0}}\|_{1}},\forall b\in[d]$. The claim follows, since $P_{k}(C)-P_{k-1}(C^{\prime})=\sum_{b}(p_{b}-p_{b}^{\prime})\leq\sum_{b}\frac{|c_{i_{0}}^{(b)}-c_{j_{0}}^{(b)}|}{\|c_{i_{0}}-c_{j_{0}}\|_{1}}=1$.
If ${c^{\prime}_{u}}^{(b)}\neq{c^{\prime}_{v}}^{(b)}$,
${p_{b}}^{\prime}\geq\frac{|{c^{\prime}_{u}}^{(b)}-{c^{\prime}_{v}}^{(b)}|}{\|c^{\prime}_{u}-c^{\prime}_{v}\|_{1}}\geq\frac{|{c_{u}^{(b)}}-{c_{v}^{(b)}}|-|{c_{i_{0}}^{(b)}}-{c_{j_{0}}^{(b)}}|}{\|c_{u}-c_{v}\|_{1}}.$
The second inequality follows from 5.13. Thus
$p_{b}-p_{b}^{\prime}\leq\frac{|{c_{i_{0}}^{(b)}}-{c_{j_{0}}^{(b)}}|}{\|c_{u}-c_{v}\|_{1}}\leq\frac{|{c_{i_{0}}^{(b)}}-{c_{j_{0}}^{(b)}}|}{\|c_{i_{0}}-c_{j_{0}}\|_{1}}.$
If ${c^{\prime}_{u}}^{(b)}={c^{\prime}_{v}}^{(b)}$, it means $|{c_{u}^{(b)}}-{c_{v}^{(b)}}|\leq|{c_{i_{0}}^{(b)}}-{c_{j_{0}}^{(b)}}|$, thus
$p_{b}-p_{b}^{\prime}\leq p_{b}=\frac{|c_{u}^{(b)}-c_{v}^{(b)}|}{\|c_{u}-c_{v}\|_{1}}\leq\frac{|{c_{i_{0}}^{(b)}}-{c_{j_{0}}^{(b)}}|}{\|c_{i_{0}}-c_{j_{0}}\|_{1}}.$
Solving the recurrence given by 5.14 with the boundary condition that for a set $C^{\prime\prime}$ of two points, $P_{2}(C^{\prime\prime})=1$, we get that $P_{n}(C)\leq n$. Since $P_{n}(C)$, for the set of $n$ points, is by definition, equal to $\sum_{b\in[d]}p^{(b)}$, the statement of the lemma follows. ∎
Claim 5.15.
With probability $1-\delta^{\prime}$, we have $|I|\leq 2Tn$.
Proof.
Using Chernoff bound, we have
$\Pr\left[|I|-\mathbb{E}\left[|I|\right]\geq\mathbb{E}\left[|I|\right]\right]\leq\exp(-\mathbb{E}[|I|]/3)\leq\exp(-T/3)\leq\delta/4.$
Thus, with probability $41-\delta/$ 4, we have $|I|\leq 2E[|I|]\leq 2Tn$. ∎
Lemma 5.16.
With probability $1-\delta^{\prime}$, the space complexity of Algorithm 3 is $O(\frac{n\log^{2}(n/\delta)\log(d)}{\epsilon^{5}\delta^{4}})$ wordsize, and the number of coordinates that we query is $O(\frac{n\log(n/\delta)}{\epsilon^{3}\delta^{2}})$.
Proof.
By 5.15, we have that the number of coordinates of $q$ that we query is $|I|=O(Tn)=O(\frac{n\log(n/\delta)}{\epsilon^{3}\delta^{2}})$. Next we analyze the space complexity of the data structure. Matrix $M$ is of wordsize $O(m|I|)$, $M(R)$ is of wordsize $O(mn)$, $I$ is of wordsize $O(|I|\log(d))$. Thus the space complexity of the our data strcuture is $O(m|I|+mn+|I|\log(d))=O(\frac{n\log^{2}(n/\delta)\log(d)}{\epsilon^{5}\delta^{4}})$ wordsize. ∎
@@ 5.1.4 Space and Query Complexity
5.2 Data Structure for $\ell_{2}$-metric with Near-linear Space and Query Time

### v2

@@ B.1 Missing Proofs from Section 4 for $\ell_{1}$-metric
$\mathbb{E}[\|r_{i}-u\|_{1}]$ $=\mathbb{E}\left[\sum_{b\in I}\left|r_{i}^{(b)}-u^{(b)}\right|\right]$
$=\mathbb{E}\left[\sum_{t\in[T],b\in[d]}\mathbbm{1}_{\{\text{in iteration $t$, we add $b$ to $I$\}}}\left|r_{i}^{(b)}-u^{(b)}\right|\right]$
$=\sum_{t\in[T],b\in[d]}p^{(b)}\left|r_{i}^{(b)}-u^{(b)}\right|=\sum_{t\in[T],b\in[d]}p^{(b)}\left|\frac{c_{i}^{(b)}}{p^{(b)}}-\frac{q^{(b)}}{p^{(b)}}\right|$
$=\sum_{t\in[T],b\in[d]}\left|c_{i}^{(b)}-q^{(b)}\right|=T\|c_{i}-q\|_{1}.$
@@
$\|r_{i^{*}}-u\|_{1}\leq\frac{1}{\delta^{\prime}}\mathbb{E}[\|r_{i}-u\|_{1}]\leq\frac{T}{\delta^{\prime}}\|c_{i^{*}}-q\|_{1}.$
Proof B.2 (Proof of Lemma 4.5).
@@
$\|b-q\|_{1}+\|a-q\|_{1}\geq\mathsf{dist}_{\mathsf{Bad}(a,b,q)}(b,q)+\mathsf{dist}_{\mathsf{Bad}(a,b,q)}(a,q)$
$\geq\sum_{z\in\mathsf{Bad}(a,b,q)}\left(\frac{8}{\epsilon\delta^{\prime}}+\frac{8}{\epsilon\delta^{\prime}}\right)|b^{(z)}-a^{(z)}|.$
@@
$\frac{\mathsf{dist}_{\mathsf{Good}(a,b,q)}(b,a)}{\|b-a\|_{1}}\geq\frac{14-\epsilon\delta^{\prime}}{16}\geq\frac{13}{16}.$
Proof B.3 (Proof of Lemma 4.6).
@@
Notice $\forall t\in[T],\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]=\mathsf{dist}_{\mathsf{Good}}(c_{i},q)$ and $\mathbb{E}[\sum_{z\in\mathsf{Good}}Q_{t,z}]=\mathsf{dist}_{\mathsf{Good}}(c_{j},q)$. Because
$\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]$ $=\sum_{z\in\mathsf{Good}}\mathbb{E}[P_{t,z}]=\sum_{z\in\mathsf{Good}}p^{(z)}\cdot\frac{|q^{(z)}-c_{i}^{(z)}|}{p^{(z)}}=\sum_{z\in\mathsf{Good}}|q^{(z)}-c_{i}^{(z)}|$
$=\mathsf{dist}_{\mathsf{Good}}(c_{i},q).$
@@
$\Pr\left[\left|\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}-\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right|\geq\epsilon_{2}\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right]$
$\leq$ $2\exp\left(-\frac{\epsilon_{2}^{2}T\mathbb{E}[\sum_{z\in[d]}P_{t,z}]}{3(1+\frac{8}{\epsilon\delta^{\prime}})\|c_{i}-c_{j}\|_{1}}\right)$
$=$ $2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{8}{\epsilon\delta^{\prime}})}\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},q)}{\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{j})}\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{j})}{\|c_{i}-c_{j}\|_{1}}\right)$
$\leq$ $2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{8}{\epsilon\delta^{\prime}})}\cdot\frac{1}{2}\cdot\frac{13}{16}\right)\leq\frac{\delta^{\prime}}{2n},$
@@
$\Pr\left[\left|\sum_{t\in[T],z\in\mathsf{Good}}Q_{t,z}-\mathbb{E}[\sum_{t\in[T],z\in\mathsf{Good}}Q_{t,z}]\right|>\frac{T}{8}\cdot\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{j})\right]$
$\leq$ $2\exp\left(-\frac{T\cdot\frac{1}{8}\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{j})}{3(1+\frac{8}{\epsilon\delta^{\prime}})\|c_{i}-c_{j}\|_{1}}\right)$
$\leq$ $2\exp\left(-\frac{T}{24(1+\frac{8}{\epsilon\delta^{\prime}})}\frac{\mathsf{dist}_{\mathsf{Good}}(c_{i},c_{j})}{\|c_{i}-c_{j}\|_{1}}\right)\leq\frac{\delta^{\prime}}{2n}.$
@@
The third item of the lemma is proved.
Proof B.4 (Proof of Lemma 4.7).
We know that $\|r_{i^{*}}-u\|_{1}\leq\frac{T}{\delta^{\prime}}\|c_{i^{*}}-q\|_{1}$ holds with probability $1-\delta^{\prime}$ by 4. Thus, it is sufficient to show the statement of the lemma holds under this assumption with probability $1-\delta^{\prime}$.
@@
Given these conditions, the following 9 gives a lower bound on the ratio $\frac{g_{1}^{\prime}+b_{1}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}$. We defer the proof of the claim to the appendix.
Claim 9.
@@
Applying 9, we get that $\frac{g_{1}^{\prime}+b_{1}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq 1+\epsilon\delta^{\prime}/4$, which implies $\frac{\|r_{i}-u\|_{1}}{\|r_{i^{*}}-u\|_{1}}\geq 1+\epsilon\delta^{\prime}/4$. By union bound, this inequality holds for all $i\in[n]$ with probability $1-\delta^{\prime}$.
Claim 10.
Proof B.5.
From $g_{1}^{\prime}\approx_{\epsilon/16}g_{1}$ and $w^{\prime}\approx_{\epsilon/16}w$, we have $g_{1}^{\prime}-w^{\prime}\geq\left(1-\frac{\epsilon}{16}\right)g_{1}-\left(1+\frac{\epsilon}{16}\right)w.$ Let $\rho=\frac{3}{4}-\frac{\epsilon}{16}$. We rewrite $\frac{g_{1}}{w}\geq 1+\frac{\epsilon}{2}$ as $\left(1-\frac{\epsilon}{16}-\rho\right)g_{1}\geq\left(1+\frac{\epsilon}{16}-\rho\right)w$. This shows $\left(1-\frac{\epsilon}{16}\right)g_{1}-\left(1+\frac{\epsilon}{16}\right)w\geq\rho(g_{1}-w)$. Thus we have $g_{1}^{\prime}-w^{\prime}\geq\rho(g_{1}-w)$.
Proof B.6 (Proof of 9).
Given $\frac{g_{1}+b_{1}}{g_{2}+b_{2}}\geq 1+\epsilon$, we start by deducting $1+\frac{\epsilon\delta^{\prime}}{8}$ on both sides: $\frac{g_{1}+b_{1}-(1+\frac{\epsilon\delta^{\prime}}{8})(g_{2}+b_{2})}{g_{2}+b_{2}}\geq 1+\epsilon-(1+\frac{\epsilon\delta^{\prime}}{8})\geq\frac{7}{8}\epsilon.$ Since $\frac{b_{1}}{b_{2}}\leq 1+\frac{\epsilon\delta^{\prime}}{8}$, we get $\frac{g_{1}-(1+\frac{\epsilon\delta^{\prime}}{8})g_{2}}{g_{2}+b_{2}}\geq\frac{7}{8}\epsilon.$ Using 10 and letting $w=(1+\frac{\epsilon}{8})g_{2}$, we have $\frac{g_{1}^{\prime}-(1+\frac{\epsilon\delta^{\prime}}{8})g_{2}^{\prime}}{g_{2}+b_{2}}\geq\frac{21-\frac{7\epsilon}{4}}{32}\epsilon\geq\frac{5\epsilon}{8}.$ Since $g_{2}^{\prime}+b_{2}^{\prime}\leq\frac{1}{\delta^{\prime}}(g_{2}+b_{2})$, we get $\frac{g_{1}^{\prime}-(1+\frac{\epsilon\delta^{\prime}}{8})g_{2}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq\frac{5\epsilon\delta^{\prime}}{8}.$ Rewriting the left-hand side as $\frac{g_{1}^{\prime}-g_{2}^{\prime}+b_{1}^{\prime}-b_{2}^{\prime}-\frac{\epsilon\delta^{\prime}}{8}(g_{2}^{\prime}+b_{2}^{\prime})-b_{1}^{\prime}+(1+\frac{\epsilon\delta^{\prime}}{8})b_{2}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq\frac{5\epsilon\delta^{\prime}}{8}.$ Rearranging terms gives $\frac{g_{1}^{\prime}-g_{2}^{\prime}+b_{1}^{\prime}-b_{2}^{\prime}-b_{1}^{\prime}+(1+\frac{\epsilon\delta^{\prime}}{8})b_{2}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq\frac{5\epsilon\delta^{\prime}}{8}+\frac{\epsilon\delta^{\prime}}{8}.$ Using $\frac{b_{1}^{\prime}}{b_{2}^{\prime}}\geq 1-\frac{\epsilon\delta^{\prime}}{8}$, we get $\frac{g_{1}^{\prime}-g_{2}^{\prime}+b_{1}^{\prime}-b_{2}^{\prime}+(\frac{\epsilon\delta^{\prime}}{4})b_{2}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}=\frac{g_{1}^{\prime}-g_{2}^{\prime}+b_{1}^{\prime}-b_{2}^{\prime}-(1-\frac{\epsilon\delta^{\prime}}{8})b_{2}^{\prime}+(1+\frac{\epsilon\delta^{\prime}}{8})b_{2}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq\frac{3\epsilon\delta^{\prime}}{4}.$
@@
B.2 Data Structure for $\ell_{2}$-metric with Near-linear Space and Query Time